Find the common factors of the given terms.
 2x, 3x^2 , 4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Factors of 2x, 3x^2 and 4
2x = 2 × x
3x^2= 3 × x × x
4 = 2x2
Common factors of 2x, 3x2and 4 is 1.

Thanks